3 - Сс У (х -

y = ५ (x ५- 1)
Answer #1

Given differential equation is dy = y(102 -1) 4 - +y=x* 1 dy 1 + = X (1) 4 mate = 2, then 1 dz 3 dx 3 dy 1 dy Put * dx vt dx